OpenStudy (anonymous):

Can someone please help me! I will fan+medal Find the sum of the first six terms of the geometric sequence for which a2=0.7 and a3=0.49

OpenStudy (bibby):

what is the value of r?

OpenStudy (anonymous):

0.7

OpenStudy (bibby):

so what is the value of a1, a4, a5, a6?

OpenStudy (bibby):

well we don't even need all of the terms

OpenStudy (anonymous):

a1=0.1 a4=3.43 a5=24.01 a6=168.07

OpenStudy (bibby):

you're multiplying by 0.7 not 7. the numbers should be getting smaller after a3. im trying to find that one formula

OpenStudy (anonymous):

but to get to 0.7 you must multiply 0.1 times 7 and to get to 0.49 you must multiply 0.7 to 0.7

OpenStudy (anonymous):

if you see 0.7 times 0.7=0.49

OpenStudy (bibby):

ok I found it. \(\huge S_n=a_1(\frac{{1-r}^n}{1-r})\)

OpenStudy (bibby):

ok first thing's first, to get a1, we have to divide a2 by 0.7 because we're working backwards.

OpenStudy (anonymous):

mhmm

OpenStudy (anonymous):

thats 0.1

OpenStudy (bibby):

if a1 = 0.1 and the ratio = 0.7 a2 would be 0.07 if a1 = 1 and the ratio = 0.7 a2 would be 0.7

OpenStudy (anonymous):

oh okay

OpenStudy (bibby):

For reference, 0.7/0.7 = \(a_1\) = 1

OpenStudy (anonymous):

yeah makes sense

OpenStudy (bibby):

so where were we \(\huge S_n=a_1(\frac{{1-r}^n}{1-r})\) \(\huge S_n=1(\frac{{1-0.7}^6}{1-0.7})=\frac{0.3^6}{0.3}\)

OpenStudy (anonymous):

umm i got 0.00243

OpenStudy (bibby):

same I don't know if that makes sense

OpenStudy (bibby):

1+0.7+0.49+... should be at least 1

OpenStudy (anonymous):

but the way i did it before, the way that you thought was wrong i got 196.8

OpenStudy (bibby):

That's definitely wrong. 1, 0.7, 0.49, 0.343, 0.2401, 0.16807 is the actual series assuming r = 0.7

OpenStudy (anonymous):

hmm so lost

OpenStudy (bibby):

here, if \(\large a_2=0.7 \) \(\large a_3=0.49 \) That means the common ratio is 0.7. That also means that \(\large a_1=1\) and \(\large a_4=0.343\)

OpenStudy (anonymous):

so whats the answer?

OpenStudy (bibby):

1 + 0.7 + 0.49 + 0.343 + 0.2401 + 0.16807

OpenStudy (bibby):

I'll try it again with one of the formulas in a sec

OpenStudy (bibby):

oh wow I'm an absolute idiot

OpenStudy (anonymous):

?

OpenStudy (bibby):

\(\huge S_n=a_1(\frac{{1-(r}^n)}{1-r})\)\(\large r=0.7, n =6\) \[\large S_n=(\frac{{1-(r}^n)}{1-r})=(\frac{{1-(0.7}^6)}{1-0.7})=\frac{1-0.117649}{0.3}= \frac{0.882351}{0.3}=2.94117\]

OpenStudy (bibby):

2.94117

OpenStudy (anonymous):

thats the final answer?

OpenStudy (bibby):

yeah. I did the math too, google agrees 1 + 0.7 + 0.49 + 0.343 + 0.2401 + 0.16807 =2.94117

OpenStudy (anonymous):

thanks so much!

OpenStudy (bibby):

sorry about getting caught up with something as stupid as an arithmetic mistake

OpenStudy (anonymous):

its geometric :) and its okay

OpenStudy (bibby):

I can't tell if that's a joke or not :p I meant the fact that I subtracted before raising the power. PEDMAS